# Penrose’s eight-conic theorem via Penrose’s eight-quadric theorem

### v3

@@ 2.1.1 Double contact
when $P$ is incident to $p$. In particular, the intersection points of $B$ with $A$ are always the intersection points of $p$ with $A$ and the tangent lines there are always the same, independently of $c$. Thus $A$ and $B$ are said to be in double contact with chord of contact $p$ and the one parameter family of conics defined by equation (6) is called a double contact pencil. The property of being in double contact is clearly self dual: two point-wise conics share tangent lines at two common points if and only if the two corresponding linewise conics share turning points at two common lines, and the fact that equation (6) is self dual, i.e., that the linewise conics of a double contact pencil can be expressed as the linear span of a regular conic and a double point: the point where the shared tangent lines meet, can be checked by taking cofactor matrices. We call this double point the apex of contact. Therefore the polarity property expressed by equation (7) can be stated as
@@
The polarities on two conics in double contact act the same when restricted to a mapping between the points of the chord of contact and the lines of the apex of contact.
@@ 2.2 Quadrics and projective space
Similar definitions hold in projective 3 space $\mathbb{R}P^{3}$, using a 4 dimensional vector space $\mathtt{V}$ where now the dual space ${\mathbb{R}P^{3}}^{*}$ is the space of planes. A correlation is defined by the property that 3 collinear points map to 3 coaxial planes (three planes going through the same line: we use the word "axis" to emphasize the dual role that a line plays as the intersection of two planes in contrast to its more familiar role of connecting two points). A polarity is defined the same way but we now refer to the surface of points incident to their polar planes as a point-wise quadric. There are now 3 different cases to consider:
  • If the kernel $\mathrm{ker}A$ has dimension $1$, it represents the vertex of a cone which constitutes the points of the quadric.
@@
  • If the kernel $\mathrm{ker}A$ has dimension $3$, it is a (doubled) plane. If this plane is called $p$ then
@@
  • If the kernel $\mathrm{ker}A$ has dimension $2$, it represents the connecting line of a point pair. If these points are called $P$ and $Q$ then,
@@
  • If the kernel $\mathrm{ker}A$ has dimension $3$, it is a (doubled) point. If this point is called $P$ then
@@
The set of geometric elements represented by vectors in the kernel of $A$ is known as the degenerate set, for point-wise quadrics: the vertex of a cone, the points on the axis of a plane pair, the points on a double plane, and for plane-wise quadrics: the plane of a conic, the planes through the connecting line of a point pair or the planes through a double point. For degenerate point-wise quadrics, the polarity mapping is only defined from points, not contained in the degenerate set, to planes and vice versa for plane-wise quadrics. For point-wise quadrics:
@@
2.2.1 Notation
We write the point-wise, resp., plane-wise aspect of a quadric as $\mathbf{{Q}}$, resp., $\mathbf{{Q}}^{*}$, and the pair $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ for the complete quadric. When either is regular, then so is the other and the relationship is given by $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda\neq 0$. Otherwise, both are degenerate and they satisfy the relationship $\mathbf{{Q}}\mathbf{{Q}}^{*}=0$. In discussions limited to regular quadrics, we mostly refer to a quadric simply in its point-wise form as $\mathbf{{Q}}$, since $\mathbf{{Q}}^{*}$ is uniquely determined. When degenerate quadrics play a role we employ the full notation introduced here.
@@ 2.2.2 Ring contact
meet in the conic where $p$ intersects $A$ and they have the same tangent planes at all these shared points. We call $p$ the ring plane and
refer to its polar point as the ring points. By the same reasoning as in the conic case, we have
@@
In the case where ring point is incident to the ring plane, i.e. the ring plane is a tangent plane and the ring point is its point of contact, the conic of ring contact is a (doubled) pair of lines. This case can be distinguished as X-contact though we include it under the umbrella term “ring contact”.
@@ 2.2.3 Some contact lemmas
We conclude this section with a few lemmas that will be useful for the proof the 8-quadric theorem.
@@
If $A,B$ and $C$ are in ring contact then there exist planes $p,q$ and $r$ such that
@@
On the other hand, if the chords of contact $p,q,r$ all meet in a point, then the span of $p^{2},q^{2},r^{2}$ contains only line pairs through this common point so we have the following lemma:
@@
If 3 conics are mutually in double contact with concurrent chords of contact, then all 3 are in the same double contact pencil.
@@ 3.1 Notation and terminology
Next, we introduce notation for the vertices and faces of the cube. Consult Fig. 5, left. We decompose the cube vertices into two interlocked tetrahedra, and label the first group of vertices $\mathbf{{S}}_{i},i\in(0,1,2,3)$ and the second $\mathbf{{T}}^{i}$, where $\mathbf{{T}}^{i}$ is opposite to $\mathbf{{S}}_{i}$ with respect to the center of the cube. Write the face of the cube with corners $\mathbf{{T}}^{i}\mathbf{{S}}_{k}\mathbf{{T}}^{j}\mathbf{{S}}_{m}$ as $F^{ij}_{km}$ where $ijkm$ is a permutation of $0123$.
In condition (ii) of the 8-conic theorem, we call the common point of the chords of contact on a face the face point of the face; for condition (ii) of the 8-quadric theorem we call the common axis of the ring planes the face axis of the face.
Furthermore, in what follows we assume that the empty vertex of a Penrose 7-configuration is $\mathbf{{T}}^{0}$.
@@
The face figure of a Penrose conic configuration. Consider two conics corresponding to opposite nodes in a face of the cube, e.g. $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$. Then by (6) they can be written as $\mathbf{{T}}^{1}=\mathbf{{S}}_{0}+pp^{T}$ and $\mathbf{{T}}^{3}=\mathbf{{S}}_{0}+qq^{T}$ with $p,q$ being the chords of contact between $\mathbf{{S}}_{0}$ and $\mathbf{{T}}^{i}$. Hence $\mathbf{{T}}^{1}-\mathbf{{T}}^{3}=pp^{T}-qq^{T}=\frac{1}{2}(de^{T}+ed^{T})$ with $d=p+q$ and $e=p-q$ being a line pair of common chords in the pencil spanned by $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$. Thus the chords of contact meet at the point of intersection of a pair of common chords $e,d$. Since there are at most 3 such pairs of common chords the set of conics being in double contact with $\mathbf{{T}}^{1},\mathbf{{T}}^{3}$ splits into at most 3 families, such that condition (ii) holds automatically for all the conics $\mathbf{{S}}_{0},\mathbf{{S}}_{2}$ in the same family. All these conics can explicitly be written in terms of $d,e$ and $\mathbf{{S}}_{0}$ (see Figure 5 and [Sal17, p.262]).
@@ 3.2 Penrose’s 8-conic theorem
The extrusion needed for the 7-configurations encountered in the following proofs is provided by the following lemma, whose proof appears below in Sec. 4.
@@ 3.3 Penrose’s 8-quadric theorem
These will both be nonzero since otherwise, if we had $P^{n}_{i}=P^{n}_{j}$ or $p^{n}_{i}=p^{n}_{j}$, $\mathbf{{S}}_{i}$ and $\mathbf{{S}}_{j}$ would have to be in ring contact along the common conic of intersection with $\mathbf{{T}}^{n}$, contradicting Lemma 3.2.
@@
For a regular Penrose 7-quadric configuration: If two of the face axes at $\mathbf{{S}}_{0}$ are distinct, then at each complete vertex, all three face axes are distinct. In this case, all ring planes and face axes are concurrent in a point $O$. Otherwise there is a single face axis $\Omega$ common to all ring planes. Dually, all ring points and face spears are concurrent in a plane $o$ or there is a single face spear $\omega$ common to all ring points. $O$ (resp., $\Omega$) and $o$ (resp., $\omega$) are a polar pair with respect to all quadrics of the configuration.
@@
This contradicts the assumption that two of the three face axes $\Lambda^{ij}_{0k}$ are distinct. Hence all three are distinct if any two are.
@@
Hence the three axes at $\mathbf{{S}}_{0}$ are distinct and do not lie in a plane, and the ring planes are therefore distinct, and have a common point $O$, which is therefore also a common point of the axes.
Apply the same argument to the other complete vertices $\mathbf{{T}}^{i}$ to conclude that the axes of the faces meeting at $\mathbf{{T}}^{i}$ have a common point $O^{\prime}$. We show that $O^{\prime}=O$. Indeed, two of these axes belong to $\{\Lambda^{ij}_{0k}\}$, that were shown above to meet in $O$. So $O^{\prime}=O$. Since this applies to all 3 incomplete faces, all 6 axes are concurrent in $O$.
We now consider the case that two of the axes at $\mathbf{{S}}_{0}$ are the same. Then by the above, all three axes $\Lambda^{ij}_{0k}$ are the same, call it $\Omega$. By reasoning similar to the above, we can propagate this condition to the three remaining face axes, to show they also have to be equal to $\Omega$, and consequently that all ring planes pass through $\Omega$.
@@
We will also need this standard geometric result (see e.g. [Zie12, §4.3]):
@@
Note that the conics cannot lie in a plane since two such conics have 4 points in common. The reader is also referred to the “coordinate version” of the 8-quadric proof below, which provides an independent proof of the lemma.
@@ 3.3.1 The proof
We also assume for the moment that $o$ is not incident to $O$. This is equivalent to assuming the 7 quadrics of the configuration do not all share a common contact element. We postpone to Sec. 4.1 the proof that this condition is not necessary, since it depends on elements of the proof of Lemma 3.1.
Figure 6 depicts the completed cube configuration which we must construct. There are 6 axes corresponding to condition (ii), one for each face. Each face axis is the common line of the 4 ring planes on the edges of the face. As mentioned above, we restrict our attention to the first case in Lemma 3.3, whereby the 3 axes of the complete faces of the 7-configuration are not all identical and have a common point $O$, as shown in Figure 6. The dual plane $o$ is taken as the plane at infinity in this figure, thus the cones of ring contact are all cylinders through the conics $\mathbf{{R}}^{k}_{j}$ where $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{k}$ make ring contact. The 4 ring points on a face lie on six distinct spears in $o$.
@@
We now show condition (i) is also fulfilled. The polarity of $\mathbf{{R}}^{0}_{i}$ in the plane $p^{0}_{i}$ maps a point $P$ to the line of intersection of $p^{0}_{i}$ and the polar plane of $P$ with respect to any quadric containing $\mathbf{{R}}^{0}_{i}$: $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{0}$ in particular. By equations (10) and (11), and Lemma 2.2, the polarities on $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$ and $\mathbf{{T}}^{0}$ all behave the same way when restricted to points on $\Lambda^{0n}_{jk}$ and planes through $\lambda^{0n}_{jk}$. Combining these two facts and applying them to the cases $i=j$ and $i=k$, any point on $\Lambda^{0n}_{jk}$ is mapped under the respective polarities of $\mathbf{{R}}^{0}_{j}$ and $\mathbf{{R}}^{0}_{k}$ to two lines which together span the polar plane of $P$ with respect to any of the quadrics $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$, $\mathbf{{T}}^{0}$. Finally, by considering the same scenario with $\Lambda^{0k}_{jn}$ in place of $\Lambda^{0n}_{jk}$, and appealing to the definitions of $P^{0}_{j}$ and $p^{0}_{j}$ in equations (12) and (13), we conclude that the polarities on $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{0}$ agree when restricted to planes through $P^{0}_{j}$ and points on $p^{0}_{j}$ so $\mathbf{{T}}^{0}$ and $\mathbf{{S}}_{j}$ are in ring contact by Lemma 2.2. ∎

### v4

@@ 2.1.2 Double contact
when $P$ is incident to $p$. In particular, the intersection points of $B$ with $A$ are always the intersection points of $p$ with $A$ and the tangent lines there are always the same, independently of $c$. Thus, $A$ and $B$ are said to be in double contact with the chord of contact of the line of contact $p$, and the pencil of conics defined by Eq. (6) is called a double contact pencil. The property of being in double contact is clearly self-dual. Indeed, two point-wise conics share tangent lines at two common points if and only if the two corresponding line-wise conics share turning points at two common lines. Furthermore, Eq. (6) is self-dual, that is, the line-wise conics of a double contact pencil can be expressed as the linear span of a regular conic and a double point, given by the meet of the two lines, as can be checked by taking cofactor matrices. We call this double point the point of contact. Therefore, the polarity property expressed by Eq. (7) can be stated as
@@
The polarities on two conics in double contact act the same when restricted to a mapping between the points of the chord of contact and the lines of the point of contact.
@@ 2.2 Quadrics and projective space
Similar definitions hold in projective 3-space $\mathbb{R}P^{3}$, using a 4-dimensional vector space $\mathtt{V}$ where now the dual space ${\mathbb{R}P^{3}}^{*}$ is the space of planes. A correlation is defined by the property that 3 collinear points map to 3 coaxial planes (three planes going through the same line: we use the word "axis" to emphasize the dual role that a line plays as the intersection of two planes in contrast to its more familiar role of connecting two points, for which we use the term “spear”. A polarity is defined the same way, but we now refer to the surface of points incident to their polar planes as a point-wise quadric. There are now 3 different cases to consider:
  • If the kernel $\mathrm{ker}A$ has dimension $1$, it represents the vertex of a cone that constitutes the points of the quadric.
@@
  • If the kernel $\mathrm{ker}A$ has dimension $3$, it is a (double) plane. If this plane is called $p$ then
@@
  • If the kernel $\mathrm{ker}A$ has dimension $2$, it represents the joining line of a point pair. If these points are called $P$ and $Q$ then
@@
  • If the kernel $\mathrm{ker}A$ has dimension $3$, it is a (double) point. If this point is called $P$ then
@@
The set of geometric elements represented by vectors in the kernel of $A$ is known as the degenerate set. For point-wise quadrics this is: the vertex of a cone, the points of the axis of a plane pair, the points of a double plane; and for plane-wise quadrics: the plane of a conic, the planes of the joining line of a point pair, the planes of a double point. For degenerate point-wise quadrics, the polarity mapping is only defined for points not in the degenerate set to planes, and vice versa for plane-wise quadrics. For point-wise quadrics:
@@ 2.2.1 Ring contact
meet in the conic where $p$ intersects $A$ and they have the same tangent planes at all these shared points. We call $p$ the ring plane and refer to its polar point as the ring points. By the same reasoning as in the conic case, we have
@@
In the case where the ring point is incident to the ring plane, that is, the ring plane is a tangent plane and the ring point is its point of contact, the conic of the ring contact is a (double) pair of lines. This case can be distinguished as X-contact although we include it under the umbrella term “ring contact”.
@@ 2.2.2 Some contact lemmas
We conclude this section with a few lemmas that will be useful to prove the 8-quadric theorem. By ring contact triangle we mean a set of 3 ring contact pencils that pairwise have one conic in common.
@@
If $A,B$ and $C$ are in ring contact, then there exist planes $p,q$ and $r$ such that
@@
On the other hand, if the chords of contact $p,q,r$ all meet in a point, then the span of $p^{2},q^{2},r^{2}$ contains only line pairs through this common point, so we have the following lemma:
@@
If 3 conics are pairwise in double contact with concurrent chords of contact, then all 3 are in the same double contact pencil.
The following extended notation will be used below in Sec. 3.3.
2.3 Notation for complete quadrics
Every quadric has a point-wise, or primary, aspect as well as a plane-wise, or dual, aspect. We write the primary (dual) form of a quadric as $\mathbf{{Q}}$ ($\mathbf{{Q}}^{*}$). When does the pair $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ represent a quadric? When either is regular, then so is the other, and the relationship is given by $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda\neq 0$. Otherwise, both are degenerate and satisfy the relationship $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda=0$, which can be seen by taking limits of regular quadrics. Then a necessary condition for $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ to represent a quadric is $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$. If in addition both are not zero, we call the pair a complete quadric. In discussions limited to regular quadrics, we usually refer to a quadric simply in its point-wise form as $\mathbf{{Q}}$, since $\mathbf{{Q}}^{*}$ is uniquely determined. When degenerate quadrics play a role, we employ the full notation introduced here.
All the above also applies to conics, if plane-wise is replaced by line-wise.
@@ 3.1 Notation and terminology
Next, we introduce notation for the vertices and faces of the cube. Consult Fig. 5, left. We decompose the cube vertices into two interlocked tetrahedra and label the first group of vertices $\mathbf{{S}}_{i},i\in(0,1,2,3)$ and the second $\mathbf{{T}}^{i}$, where $\mathbf{{T}}^{i}$ is opposite to $\mathbf{{S}}_{i}$ with respect to the center of the cube. Write the face of the cube with corners $\mathbf{{T}}^{i}\mathbf{{S}}_{k}\mathbf{{T}}^{j}\mathbf{{S}}_{m}$ as $F^{ij}_{km}$ where $ijkm$ is a permutation of $0123$.
In condition (ii) of the 8-conic theorem, we call the common point of the chords of contact on a face the face point of the face; for condition (ii) of the 8-quadric theorem, we call the common axis of the ring planes the face axis of the face.
Furthermore, in what follows, we assume that the empty vertex of a Penrose 7-configuration is $\mathbf{{T}}^{0}$.
@@
3.2 The face figure of a Penrose conic configuration
In Fig. 5 (left), consider two conics corresponding to opposite nodes in a face of the cube, for example, $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$. Then by (6) they can be written as $\mathbf{{T}}^{1}=\mathbf{{S}}_{0}+pp^{T}$ and $\mathbf{{T}}^{3}=\mathbf{{S}}_{0}+qq^{T}$ with $p,q$ being the chords of contact between $\mathbf{{S}}_{0}$ and $\mathbf{{T}}^{i}$. Hence $\mathbf{{T}}^{1}-\mathbf{{T}}^{3}=pp^{T}-qq^{T}=\frac{1}{2}(de^{T}+ed^{T})$ with $d=p+q$ and $e=p-q$ being two lines in the pencil spanned by $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$. Thus, the chords of contact meet at the carrying point of the line pair $ed$. By Lemma 2.1, there are at most 3 line pairs in a regular pencil of conics. Hence, the set of conics that are in double contact with $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$ splits into at most 3 families such that condition (ii) of the 8-conic theorem holds automatically for all conics $\mathbf{{S}}_{0},\mathbf{{S}}_{2}$ in the same family. All these conics can be explicitly written in terms of $d,e$ and $\mathbf{{S}}_{0}$ (see [Sal17, p.262]).
@@ 3.3 Penrose’s 8-conic theorem
The extrusion needed for the 7-configurations encountered in the following proofs is provided by the following lemma, whose proof appears below in Sect. 4.
@@ 3.4 Penrose’s 8-quadric theorem
These will both be non-zero since otherwise, if we had $P^{n}_{i}=P^{n}_{j}$ or $p^{n}_{i}=p^{n}_{j}$, $\mathbf{{S}}_{i}$ and $\mathbf{{S}}_{j}$ would have to be in ring contact along the common conic of intersection with $\mathbf{{T}}^{n}$, contradicting Lemma 3.2.
@@
For a regular Penrose 7-quadric configuration: If two of the face axes at $\mathbf{{S}}_{0}$ are distinct, then at each complete vertex, all three face axes are distinct. In this case, all ring planes and face axes are concurrent in a point $O$. Otherwise, there is a single face axis $\Omega$ common to all ring planes. Dually, all ring points and face spears are concurrent in a plane $o$ or there is a single face spear $\omega$ common to all ring points. $O$ (resp., $\Omega$) and $o$ (resp., $\omega$) are a polar pair with respect to all quadrics of the configuration.
@@
This contradicts the assumption that two of the three face axes $\Lambda^{ij}_{0k}$ are distinct. Hence, all three are distinct if any two are.
@@
Hence, the three axes at $\mathbf{{S}}_{0}$ are distinct and do not lie in a plane, and the ring planes are therefore distinct, and have a common point $O$, which is therefore also a common point of the axes.
Apply the same argument to the other complete vertices $\mathbf{{T}}^{i}$ to conclude that the axes of the faces meeting at $\mathbf{{T}}^{i}$ have a common point $O^{\prime}$. We show that $O^{\prime}=O$. Indeed, two of these axes belong to $\{\Lambda^{ij}_{0k}\}$, which were shown above to meet in $O$. So $O^{\prime}=O$. Since this applies to all 3 incomplete faces, all 6 axes are concurrent in $O$.
We now consider the case where two of the axes at $\mathbf{{S}}_{0}$ are the same. Then by the above, all three axes $\Lambda^{ij}_{0k}$ are the same, call it $\Omega$. By reasoning similar to the above, we can propagate this condition to the three remaining face axes, to show that they also have to be equal to $\Omega$, and consequently that all ring planes pass through $\Omega$.
@@
We will also need this standard geometric result (see, e.g. [Zie12, §4.3]):
@@
Note that the conics cannot lie in a plane since two of such conics have 4 points in common. The reader is also referred to the “coordinate version” of the 8-quadric proof below, which provides an independent proof of the lemma.
@@ 3.4.1 The proof
We also assume for the moment that $o$ is not incident to $O$. This is equivalent to assuming that the 7 quadrics of the configuration do not all share a common contact element. We postpone to Sect. 4.1 the proof that this condition is not necessary, since it depends on elements of the proof of Lemma 3.1.
Figure 6 depicts the completed cube configuration that we must construct. There are 6 axes corresponding to condition (ii), one for each face. Each face axis is the common line of the 4 ring planes on the edges of the face. As mentioned above, we restrict our attention to the first case in Lemma 3.3, whereby the 3 axes of the complete faces of the 7-configuration are not all identical and have a common point $O$, as shown in Figure 6. The dual plane $o$ is taken as the plane at infinity in this figure, thus the cones of ring contact are all cylinders through the conics $\mathbf{{R}}^{k}_{j}$ where $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{k}$ make ring contact. The 4 ring points on a face lie on six distinct spears in $o$.
@@
We now show that condition (i) is also fulfilled. The polarity of $\mathbf{{R}}^{0}_{i}$ in the plane $p^{0}_{i}$ maps a point $P$ to the line of intersection of $p^{0}_{i}$ and the polar plane of $P$ with respect to any quadric containing $\mathbf{{R}}^{0}_{i}$: $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{0}$ in particular. By equations (10) and (11), and Lemma 2.3, the polarities on $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$ and $\mathbf{{T}}^{0}$ all behave the same way when restricted to points on $\Lambda^{0n}_{jk}$ and planes through $\lambda^{0n}_{jk}$. Combining these two facts and applying them to the cases $i=j$ and $i=k$, any point on $\Lambda^{0n}_{jk}$ is mapped under the respective polarities of $\mathbf{{R}}^{0}_{j}$ and $\mathbf{{R}}^{0}_{k}$ to two lines which together span the polar plane of $P$ with respect to any of the quadrics $\mathbf{{S}}_{j}$, $\mathbf{{S}}_{k}$, $\mathbf{{T}}^{0}$. Finally, considering the same scenario with $\Lambda^{0k}_{jn}$ in place of $\Lambda^{0n}_{jk}$, and appealing to the definitions of $P^{0}_{j}$ and $p^{0}_{j}$ in equations (12) and (13), we conclude that the polarities on $\mathbf{{S}}_{j}$ and $\mathbf{{T}}^{0}$ agree when restricted to planes through $P^{0}_{j}$ and points on $p^{0}_{j}$ so $\mathbf{{T}}^{0}$ and $\mathbf{{S}}_{j}$ are in ring contact by Lemma 2.3. ∎